OpenStudy (anonymous):

find the quotient 5 7/8 divided 6 1/4 simplify

OpenStudy (lgbasallote):

\[\huge \frac{5 \frac 78}{ 6 \frac 14} \implies \frac{\frac{47}{8}}{\frac{25}{4}}\] does that help? or do you need more help?